# Moderate Nesting and Cross-Equatorial Asymmetry of Active Regions in Solar Cycle 24

### v1

@@ 5 Conclusions
Although the raw distribution of active regions appears predominantly symmetric — with 59% of longitude–time bins co-occupied by both hemispheres — Monte Carlo tests of 10,000 randomized trials show that the the observed configuration of nests is more asymmetric than 92% of the null cases.dikpati2003 found that antisymmetric modes dominate (see their figure 6)for the dynamo-generated, solar-like toroidal bands at or below latitude, i.e., when the solar activity peaks and/or starts declining. This theoretical finding may be consistent with the observed antisymmetric-type distribution of active regions.
@@
Note that solar rotation rates have been reported to depend on the magnetic activity of the Sun (hathaway1990; wang2022), with faster rotation rates occurring during less magnetically active cycles and different rotation rates found in the hemispheres due to varying levels of magnetism during a given cycle.
In a future publication, we will report on the prevalence of long-lived nests, longitudinal modes (, etc.), and the results from wavelet analysis. Deeper analysis of the asymmetry is warranted as it suggests a way we can ascertain the mechanism influencing nesting.

### v2

@@ 5 Conclusions
Although the raw distribution of individual ARs appears predominantly symmetric — with 58% of longitude–time bins co-occupied by an AR in both hemispheres — Monte Carlo tests of 10,000 randomized trials show that the the observed configuration of nests is more asymmetric than 92% of the null cases. Dikpati et al. (2003) found that antisymmetric modes dominate (see their Figure 6) for the dynamo-generated, solar-like toroidal bands at or below latitude, i.e., when the solar activity peaks and/or starts declining. This theoretical finding may be consistent with the observed antisymmetric-type distribution of active regions.
@@
The number of nests and amount of flux in nests tracks the solar cycle evolution but does not simply scale with overall activity level. While both the magnetic flux contained in nests and the number of nests peak near hemispheric sunspot maxima, periods of enhanced sunspot number can occur without a corresponding increase in nesting. The maximum number of nests at a given time were five in the North and six in the south. The ratio of the total amount of flux in the North and South hemispheres was 0.91.
Note that solar rotation rates have been reported to depend on the magnetic activity of the Sun (Hathaway and Wilson, 1990; Wang et al., 2022), with faster rotation rates occurring during less magnetically active cycles and different rotation rates found in the hemispheres due to varying levels of magnetism during a given cycle.
In a future publication, we intend to report on the prevalence of long-lived nests (lifetimes 6 months), longitudinal modes (, etc.), and the results from wavelet analysis. Deeper analysis of the asymmetry is warranted as it suggests a way we can ascertain the mechanism influencing nesting.